OpenStudy (anonymous):

She tutors for $30 per hour and also works as a receptionist for $10 per hour. Due to her class and study schedule, Tori is only able to work up to 20 hours per week, but must earn at least $200 per week. If t represents the number of hours Tori tutors and r represents the number of hours she works as a receptionist, which system of inequalities represents this scenario? t + r greater than or equal to 20 30t + 10r = 200 t + r less than or greater to 20 30t + 10r less than or greater to 200 t + r less than or greater to 20 30t + 10r = 200 t + r le

OpenStudy (mertsj):

None of those because if she can work at most 20 hours a week then: \[t+r \le 20\]

OpenStudy (mertsj):

And that choice is not there.

OpenStudy (anonymous):

t + r less than or greater to 20 30t + 10r greater than or equal to 200 That's the last choice, it cut it off.

OpenStudy (anonymous):

@Mertsj

OpenStudy (whpalmer4):

He gave you one of the inequalities. The other is \[30t+10r \ge 200\] (money from tutoring and receptionist duties must be at least $200) If you have an answer choice that has both of those inequalities, that's the one you want.

OpenStudy (whpalmer4):

by the way, does it really say "less than or greater to" in your answer choices?

OpenStudy (anonymous):

No, it has the symbols but it changed it for some reason.

OpenStudy (anonymous):

What do you mean by "both of those inequalities"? Which inequalities?

OpenStudy (whpalmer4):

Mertsj gave you \[t+r\le20\] That's one inequality. I gave you \[30t+10r \ge 200\] That's another inequality. Are they not showing up on your screen?

OpenStudy (anonymous):

Wait, so what's the answer?

OpenStudy (whpalmer4):

You have to find the answer. We provided you with the inequalities that represent the scenario.

OpenStudy (mertsj):

Do you not understand that I told you the correct choice is not present in your answers? The question could be posted wrong. The answers could be posted wrong. The problem could be posted wrong. In any case, the answer to the question posted is not posted.

OpenStudy (whpalmer4):

@mertsj go read his first response, and remember that there's some copy/paste lossage going on here that takes the symbols and substitutes weird text for them. I've seen this on a number of problems on OS.

OpenStudy (anonymous):

you figured it out yet ?